[figure: 9.7 cm]

INFLUENCE OF THE RATIOS OF THE IONIC RADII OF CATIONS AND ANIONS ON THE MAGNETIC AND ELECTRICAL PHASE TRANSITIONS IN A_{1−x}Cu_{x}Cr_{2}X_{4} (WHERE A = Zn, Cd; X = Se, S)

J. Krok-Kowalski and J. Warczewski

University of Silesia, Institute of Physics, ul. Uniwersytecka 4, 40–007 Katowice

Let us denote: r_{t} - ionic radius of cations in the tetrahedral positions, r_{o} - ionic radius of cations in the octahedral positions, r_{a} - ionic radius of anions creating tetrahedral and octahedral gaps. For all the compounds Cd_{1−x}Cu_{x}Cr_{2}X_{4} the ratios r_{t}/r_{a} are significantly greater than r_{o}/r_{a} , which means that these compounds are ferromagnets and either semiconductors (for x ≤ 0.2) or metals (for x ≥ 0.9). Thus, in this case one can say about the semiconductor - metal phase transition forced by the change of Cu concentration [1]. Similar effect is observed for the compounds with zinc but here the semiconductor - metal phase transition is accompanied by the magnetic AF - F phase transition in the region where r_{t}/r_{a} becomes lower than r_{o}/r_{a} [2].

[1] T. Groń, H. Duda, J. Warczewski, Phys. Rev. B, 41 (1990) 12424.

[2] J. Krok, J. Spałek, S. Juszczyk, J. Warczewski, Phys. Rev. B, 28 (1983) 6499.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : Dr. Jozef Krok-Kowalski

Address for correspondence : University of Silesia, Institute of Physics, ul. Uniwersytecka 4, 40–007 Katowice, Poland

Email address : krok@us.edu.pl